I have not started my class 11.I am having doubts when I read a class 11 portion.Can anyone explain in detail,the angular momentum(mvr)(Bohr Model)?I want a detailed explanation???

Grade:10

2 Answers

M Mitra
6 years ago
As de broglie wavelength=h/mvas 2*3.14*r=nh/mvSo L=mvr=nh/2*3.14where n=1,2,3,....And it is known as Bohr`s quantisation condition for angular momentum
M Mitra
6 years ago
Acc to de broglie hypothesis,the electron revolving is also associated with wave character .Hence a circular orbit can be taken to be stationary energy state only if it contains an integral no of de broglie wavelength that is 2*3.14*r=n*§As de broglie wavelength=h/mvas 2*3.14*r=nh/mvSo L=mvr=nh/2*3.14where n=1,2,3,....And it is known as Bohr`s quantisation condition for angular momentum